ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 109, 1999

Androulakis, G. and E. Odell Distorting mixed Tsirelson spaces 125
Baribaud, C. M. Closed geodesics on pairs of pants  339

Beals, R. See D. N. Rockmore

Begun, B. A remark on almost extensions of Lipschitz functions  151
Cho, J. T. A new class of contact Riemannian manifolds  299

Di Prisco, C. A. and S. Todorcevic A cardinal defined by a polarized partition relation

41

Dixon, M. R., M. J. Evans and H. Smith A Tits alternative for groups that are residually of bounded rank

53

Dow, A. and K. P. Hart ω^{∗} has (almost) no continuous images  29

Einstein-Matthews, S. M. and C. H. Lutterodt Growth of transcendental entire functions on algebraic varieties

253

Evans, M. J. See M. R. Dixon

Fuertes, Y. Some bounds for the number of coincidences of morphisms between closed Riemann surfaces

1

Hamachi, T. Correction to “Suborbits and group extensions of flows”, Israel Journal of Mathematics, Vol. 100, 1997, pp. 249–283

379

Hart, K. P. See A. Dow

Ilani, I. Zeta functions related to the group SL_{2}(ℤ_{p})  157
Junker, M. Completeness in Zariski groups  273

Kramer, L., K. Tent and H. Van Maldeghem Simple groups of finite Morley rank and Tits buildings

189

Le Borgne, S. Limit theorems for non-hyperbolic automorphisms of the torus

61

Lutterodt, C. H. See S. M. Einstein-Matthews

Mankiewicz, P. Compact groups of operators on proportional quotients of l_{1}^{n}

75

Müller, P. Hilbert’s irreducibility theorem for prime degree and general polynomials

319

Odell, E. See G. Androulakis

Peretz, R. Applications of induced resultants to polynomial maps  181

Rockmore, D. N., K.-S. Tan and R. Beals Deciding finiteness for matrix groups over function fields

93

Ronan, M. A. and J. Tits Twin trees II: Local structure and a universal construction

349

Smith, H. See M. R. Dixon

Tan, K.-S. See D. N. Rockmore

Tent, K. See L. Kramer

Tiep, P. H. Globally irreducible representations of SL_{3}(q) and SU_{3}(q) 225

Tits, J. See M. A. Ronan

Todorcevic, S. See C. A. Di Prisco

Van Maldeghem, H. See L. Kramer

Vershik, A. and O. Zeitouni Large deviations in the geometry of convex lattice polygons

13

Watson, T. C. An improved characterization of normal sets and some counter-examples

173

Weidner, M. Augmentation-filtration and normal distribution  117

Zeitouni, O. See A. Vershik